### v2

# Template-Based Visual Program Distillation

Abstract
For users with limited computational resources, visual programming or prompting large language models (LLMs) to generate executable code for visual tasks, like visual question answering (VQA), remains largely inaccessible. Even with techniques such as distillation, adapting visual programming to smaller models or specific datasets is still quite challenging due to high annotation costs. We propose a low-cost visual program distillation method that can be used for models with fewer than 1 billion parameters and requires no human-generated program annotations. We achieve this through synthetic data augmentation based on decoupling programs into higher-level skills, called templates, and their corresponding arguments. Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality visual programs with the added benefit of much faster inference.
Template-Based Visual Program Distillation
Michal Shlapentokh-Rothman  Yu-Xiong Wang  Derek Hoiem University of Illinois at Urbana-Champaign {michal5,yxw,dhoiem}@illinois.edu
1 Introduction
Visual programming Gupta and Kembhavi (2023); Subramanian et al. (2023); Surís et al. (2023) refers to generating programs that invoke visual models to solve tasks such as answering questions about images, typically by prompting a very large language model (LLM) like GPT Achiam et al. (2023) or Llama Touvron et al. (2023). Despite success in zero-shot settings, visual programming remains largely inaccessible due to high program generation costs, which come from the need for expensive hardware to run LLMs or using expensive proprietary APIs, as well as long inference time and high data costs for both human-generated program annotations and question/answer annotations. Previous efforts Khan et al. (2024) around generating visual programs on smaller LLMs have made some progress in addressing these concerns, but training and inference costs remain substantial.
Our goal is to create a visual programming system with two key characteristics:
  1. 1.
    Small Program Generator Programs should be generated by models with $<1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
  2. 2.
    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs.
To achieve this, our approach builds on the key observation that many visual programs share a common high-level structure – a unique aspect of visual programming. For example, the programs ‘Count the red chairs’ and ‘Count the green apples’ follow the same pattern but differ in their inputs. Inspired by this insight and recent advances in LLM distillation Hsieh et al. (2023), we propose a low-cost visual program distillation method. Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most 1000). A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples. The resulting annotated programs form a dataset used to train a small language model (SLM) to generate visual programs efficiently.
Method Task Approach Model Size Train Size Domain
VisRep Khan et al. (2024) Program Gen. Self-Training 7B 10% ($\approx$10K) Visual Prog.
Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-770M 12.5% Q/A, Reasoning
CodePlan Sun et al. (2024) Program Gen. LLM Distillation 770M 100% General Coding
Template-Based Aug. (Ours) Program Gen. LLM Distillation 770M 0.1% ($\approx$1K) Visual Prog.
Table 1: Comparison of approaches for program generation and LLM distillation. Our method enables LLM distillation for visual programming using only a small fraction of the training dataset and minimizing annotation costs.
During training, program annotations are synthetically generated through template-based augmentation. This strategy decouples high-level skills or plans, called templates, from question-specific concepts or arguments. For example, the template for ‘Count the red chairs’ is ‘find(arg1), verify property(arg2), count’ and the corresponding arguments are ‘red’ (arg1) and ‘chair’ (arg2). New question/program pairs are created by replacing these arguments in the question and program. By replacing ‘red’ with ‘green’ and ‘chairs’ with ‘apples,’ we generate both a new question and a corresponding program. Compared to directly training on the program annotations, training with template-based augmentation offers three key benefits: 1) better performance when less data is available, 2) increase in student/teacher prediction agreement, and 3) higher-quality generated programs. Our method uses a low-cost teacher model (e.g. GPT-Mini-4o) and costs only around \$1 per dataset. In addition, the distilled models achieve much faster inference speed than existing LLMs.
To validate our approach, we evaluate on frequently used visual question answering (VQA) datasets and compare with using human-generated annotations as well as non-augmented distillation. Moreover, we address an overlooked evaluation aspect by performing several types of evaluation, including human program verification, due to the noisiness of commonly used metrics.
All together, our results show that it is possible with few question/answer pairs to distill the visual program ability of LLMs into a much smaller model that does not require in-context examples.
2 Related Works
We outline the most relevant related work below. A high-level comparison between our work and others can be found in Table 1.
Visual Programming
A long line of work investigates generating and executing programs to perform visual tasks. Early approaches (Andreas et al., 2015, 2016; Hu et al., 2017; Johnson et al., 2017) generate programs and execute the programs with learned end-to-end neural modules. Based on the impressive code generation capabilities of LLMs (Bareiss et al., 2022), visual programming frameworks (Surís et al., 2023; Gupta and Kembhavi, 2023; Subramanian et al., 2023) generate programs given an API and in-context examples and execute the programs with large pre-trained vision models. Visual programming has been applied to many different domains and applications including visual question answering, video question answering, text-to-image generation, and robotics (Cho et al., 2024; Lu et al., 2024; Min et al., 2024; Liang et al., 2023). Unlike previous visual programming works, we only use small models ($<1$ billion parameters) to generate programs.
Visual programming is different from both LLM tool use and LLM code generation since it requires both basic reasoning and knowledge skills as well as basic code generation. One of the challenging parts of visual programming comes from the execution and evaluation due to underlying models. As noted in other works Khan et al. (2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect.
Tool-Based Finetuning
Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming or in the general field of tool-based LLMs through finetuning. One of the main challenges is the lack of program annotations for input/output pairs. Language modeling has a long history designing self-supervised tasks, especially in pre-training (Devlin et al., 2019; Raffel et al., 2020). Toolformer (Schick et al., 2024), uses a form of self-supervision, to create a training dataset to finetune an LLM on tool-use programs. For each question/answer pair in a pre-existing training dataset, an LLM is prompted to generate a corresponding program. If the program decreases the training loss (of the same LLM) then the program annotation is added to a new dataset. Then the same LLM is finetuned on the generated dataset. A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods such as VDebugger (Wu et al., 2024) and visual programming based finetuning methods (Khan et al., 2024) such as VisRep. LoRA (Hu et al., 2021) is frequently used when finetuning LLMs on the generated data.
In VisRep, Khan et al. (2024) use a self-training approach similar to toolformer for visual programming. Self-annotated programs are kept if the executed program returns the correct answer. While there are some similarities between VisRep and our work, the underlying setting is different. VisRep focuses on improving program generation on an existing LLM (7B) through self-training, while the goal of our work is to distill visual program generation from an LLM to a small model (<1B). Our work also does not require sampling across specific question types or correcting by hand different program annotations during training. We randomly sample the training set and do not write any program annotations or corrections.
Distillation
Knowledge distillation (Hinton et al., 2015; Buciluǎ et al., 2006), where a large teacher model annotates unlabeled data that is then used to finetune a smaller, weaker student model, is frequently used across many different applications including image recognition (Beyer et al., 2021), masked language modeling (Sanh et al., 2019) and commonsense knowledge (West et al., 2021). One of the difficulties of distillation is the need for a large number of unlabeled training examples which can be expensive to obtain. One way to compensate for this is to train the student model with a multi-task objective. The objective consists of a weighted sum of the cross-entropy on the original task and cross-entropy on a closely related part of the task such as the rationale in CoT (Hsieh et al., 2023). CodePlan (Sun et al., 2024) applies such an idea to code generation where the secondary objective is a natural language version or plan of the code, which has been shown to be effective in prompting-based work (Jiang et al., 2024).
One downside to multi-objective prompting is that both the teacher and the student models have to generate additional output. In CodePlan, during inference, the model first generates the plan and then generates the desired code creating long inference time. Such steps might be necessary for complex code generation but are unnecessary for visual programming. Instead we use a relatively simple data augmentation method based on abstractions (Yuan et al., 2024) or higher level plans of existing programs which we refer to as templates. Templates do not require any additional forward passes to create and can be augmented using simple techniques such as word replacement Tang et al. (2019).
Automatic In-Context Example Generation
Prompt-based methods are more effective when in-context examples or demonstrations (Wei et al., 2022) are included. In visual programming (Surís et al., 2023; Gupta and Kembhavi, 2023), manually written in-context examples are used to adapt an API to a particular dataset. However, such an approach can be time consuming and not feasible for a large training set of input/output pairs. A common practice is to have an LLM self-annotate examples and keep the examples that produce the correct answer Zhang et al. (2023); Stanić et al. (2024); Tao et al. (2024); SU et al. (2023). Then during inference, augmented generation (RAG) Lewis et al. (2020)can be performed to select the most relevant ones. We follow such an approach, referred to as auto-context generation by  Tao et al. (2024) to generate in-context examples for the teacher model.
Template image_patch = ImagePatch(image)
var1 = image_patch.find(‘arg_0’)
var2 = var2.classify(‘arg_1’)
var3 = image_patch.find(‘arg_2’)
var4 = var3.classify(‘arg_3’)
answer = bool_to_yesno(var2 == var4)
Are the cat and the tshirt the same color? arg_0 = cat arg_1 = color arg_2 = tshirt arg_3 = color
Is the sofa made from the same material as the chair? arg_0 = sofa arg_1 = material arg_2 = chair arg_3 = material
Is the vase the same shape as the table? arg_0 = vase arg_1 = shape arg_2 = table arg_3 = shape
Table 2: A template is a particular ordering of operations. The questions above all share the same template since they only differ in the arguments. We want to answer similar questions the same way and easily generate synthetic data.
3 Method
Preliminaries
Following the notation in ViperGPT Surís et al. (2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator, $\pi$, input query $q$ and prompt $p$ such that when executed with execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer. Prompt $p$ contains an API and dataset-specific in-context examples.
Figure 1: An overview of our augmentation method. Programs are first separated into templates and argument, new arguments are selected and plugged back into the question/program pair. Templates are created by renaming variables and removing question specific concepts. One single teacher generated question/program pair can turn into hundreds of new question/program pairs.
Given a teacher model, $\pi_{t}$, and a smaller student model ($<1$ billion parameters), $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs). We do not have any human provided annotations (question/program/answer triplets) but we do know the API used for the teacher model.
There are three main steps to our approach: teacher annotation, data augmentation and student training.
3.1 Teacher Generated Program Annotation
We use GPT-4o-Mini AI (2024) as our teacher model. Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation. Common practice is to have a human generate the examples. We follow a process known as Auto-Context Generation to automatically generate such examples given answer annotations only.
Figure 2: An example of our data augmentation approach. Both the new and old question have the same template so the template matcher output should predict the same template for both. The arguments for the new and old programs are different. But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
Auto-context generation follows a simple process similar to VisRep Khan et al. (2024):
  1. 1.
    Teacher model predicts a program using API and in-context examples (if any).
  2. 2.
    Generated program is evaluated.
  3. 3.
    If the answer returned by a generated program matches the ground truth answer, then the program is immediately added to the set of in-context examples. Otherwise the program is discarded.
For efficiency, the in-context examples are sampled based on similarity to the question (i.e. RAG) Lewis et al. (2020) once there are more than 50. We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the 50 highest scoring examples.
3.2 Data Augmentation
After the teacher generates a dataset of question/program pairs, our goal is to train a student model on these pairs only. At this point, corresponding answers and images of questions are not used. Since the dataset is small, we use data augmentation to create a greater variety of question/program pairs.
To understand the intuition behind our data augmentation method, consider the set of questions in  Table 2. All of these questions compare properties of two objects. The general structure of each program is the same with the only difference coming from the inputs to the functions. If we answer one of these questions correctly and know that the remaining questions have the same structure, then all of the remaining questions should have that same structure or should be consistent. An overview of our data augmentation method can be seen in Figure 1. We refer to this as ‘template-based augmentation.’
Templates
We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation. Templates are argument independent. For example, if the program is
image_patch = ImagePatch(image)
dog = image_patch.find(‘dog’)
answer = dog.classify(‘color’)
then the template would be
image_patch = ImagePatch(image)
var1 = image_patch.find(<arg>)
answer = var1.classify(<arg>)
Please see Table 3 for examples of questions and corresponding templates.
Templates can be considered high-level plans used for plan-based distillation Sun et al. (2024). The main advantage to templates is that they can be extracted directly from the program. There is no need to generate extra output or perform multiple forward passes for a single question.
Template Extraction
Given the program annotations from the teacher model, we extract templates and corresponding arguments from each program similar to the abstraction method of CRAFT Yuan et al. (2024). Extracting is quite simple: replace specific variable names with generic ones and put in placeholders for each argument. The variable renaming and extraction can be done in seconds with abstract syntax trees and regular expression matching. The code can be found in Appendix E.
Augmentation
We can use the decomposition of programs into templates and arguments to generate synthetic data similar to how masked language modeling is used in BERT Devlin et al. (2019). As can be seen in Table 2, for many questions and programs, the arguments appear directly in the question. Consider the example in Figure 2: “Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown. Once we find similar words for each, we can simply replace them in the sentence. Since we already know the template and the arguments, we also have the program. For GQA (Hudson and Manning, 2019), the possible word replacements are from ViperGPT Surís et al. (2023) and for arguments that are either not in the question or are equal to the entire question, we use GPT to generate similar questions. For VQAv2 Goyal et al. (2017), we use BERT Devlin et al. (2019) when an argument is a single word and BART Lewis (2019) to replace phrases. If possible, we replace arguments with arguments of the same type e.g. an attribute is replaced by an attribute.
Example Template
Is the blue car the same shape as the chair? Is leather jacket made of the same material as the shirt? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = filter_img(<arg_1>)
var3 = var2.classify(<arg_2>)
var4 = image_patch.find(<arg_3>)
var5 = var4.classify(<arg_4>)
answer =bool_to_yesno(var3 ==absent$=$ = var5)
What type of food is near the person? What is the vehicle next to the animal? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = var1.crop_position(<arg_1>)
var3 = var2.find(<arg_2>)
answer $=$ var3.classify(<arg_3>)
Is the car to the left or right of the tree? Is pot above or below the pan? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = image_patch.find(<arg_1>)
answer $=$ choose_relationship(var1,var2,<arg_2>)
Table 3: Some examples of questions and corresponding templates. Multi-colored words correspond to multiple arguments.
During training, each argument in each question has a 50% chance of being replaced. If an argument is to be replaced, we then uniformly sample among the possible replacements. In the GQA dataset, for some arguments, like objects, the number of possible replacements is quite large (e.g. greater than 1500), while for arguments like directions such as left, right, etc. the number of replacements is small (e.g. fewer than 10). Some examples of categories and more details on word replacement can be found in Section F.
Student Training
Given the augmented dataset, we perform LoRA-based finetuning on the student model. Since the augmentation method produces both questions and corresponding programs, we use next-token prediction loss for training.
4 Experiments
Experimental Setup
For all experiments, the teacher model is GPT-4o-mini, the student model is CodeT5 Wang et al. (2023) and is trained with LoRA Hu et al. (2021). We use the GQA dataset Hudson and Manning (2019) and VQAv2 Goyal et al. (2017) datsets for our experiments. We evaluate on the full GQA test-dev split and randomly sample 10000 questions from the VQAv2 validation split. GQA is evaluated using exact match and VQAv2 is evaluated based on annotator/answer agreement as in the original benchmark.
We use a slightly modified API from the original ViperGPT paper  Surís et al. (2023). The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail. Please see Appendix Section C for more details. For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al. (2023) for general visual queries, Owl-ViT2 Minderer et al. (2024) for detection and CLIP Radford et al. (2021) for classification.
Auto-Context Generation
Before training the student model, we adapt the teacher to the VQA domain. In Table 4, we investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness. As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1000 programs (yielding hundreds of validated programs).
Dataset Human Generated Auto-Context Generated Performance
GQA 0 0 35.1
25 0 38.5
0 474 (out of 1000) 43.1
25 510 (out of 1000) 43.1
VQAv2 0 0 47.4
25 0 54.2
25 283 (out of 500) 57.8
0 286 (out of 500) 60.3
Table 4: Teacher performance compared across varying numbers of human-generated and auto-context-generated in-context examples.
As shown in Table 4, human-generated programs provide significant improvement vs. no examples, but provide no further benefit given self-generated examples. This finding supports using self-generation to improve the teacher model, as question-answer pairs require much less expertise and time to provide.
Data Usage GQA VQAv2
Aug. No Aug. Aug. No Aug
0.1% 41.9 (+0.8) 41.1 61.1 (+0.3) 60.8
0.05% 39.7 (+4.1) 35.6 59.9 (+9.1) 50.8
0.02% 34.2 (+4.2) 30.0 49.0 (+7.8) 41.2
Table 5: Answer accuracy with and without augmentation across different dataset sizes: 1000 (0.1%), 500 (0.05%), and 250 (0.02%) for GQA and 500 (0.1%), 250 (0.05%) and 125 (0.02%) for VQAv2. Performance gain from augmentation is higher when using 0.05% and 0.02% of the dataset compared to 0.1%.
Augmentation vs Non-Augmentation
When training the student model, we use templates and variable substitution to create a more diverse set of programs for training. We evaluate the effect of template-based data augmentation in two ways. In Table 5, we analyze answer accuracy as a function of varying dataset sizes. The results indicate that data-augmentation increases the answer accuracy of the student model, with the largest benefits when few generated programs are available for training.
To better evaluate the effectiveness of distillation, we also evaluate whether the teacher and student provide the same answer. In Table 6, we see that augmentation leads to much higher rates in student-teacher agreement, even when the overall improvement in accuracy is small (see Table 5).
GQA VQA
With Augmentation 64.5 76.1
No Augmentation 59.8 72.0
Table 6: Student/Teacher Prediction Agreement. Data augmentation has a much larger effect on student/teacher prediction compared to answer accuracy.
Cost and Efficiency
One of our objectives is keeping annotation costs low. As shown in Table 7, the total annotation cost for program auto-generation on both datasets is less than a dollar. GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question. By using a small model, inference speed also greatly increases.
GQA VQAv2
Annotation Cost \$0.69 \$0.26
CodeT5 Inference Time (q/s) 35 32.7
Llama-3.2-1B Inference Time (q/s) 0.32 0.33
GPT-4o Mini Inference Time (q/s) 1.3 1.3
Table 7: Annotation cost and inference time for several different models. Our method has low cost and high inference speed with CodeT5.
Comparison with Related Work and Program Evaluation
We compare our approach with the most closely related setting: prompt-based visual program generation. Other works, such as VisRep Khan et al. (2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons. Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
Method GQA VQA
Answer Acc Program Acc Answer Acc Program Acc
GPT-4o-Mini (w/Human Examples) 38.5 78 54.2 92
Llama 3.2 1B (w/Human Examples) 19.0 22 23.3 40
Code-T5 Distilled w/o Augmentation (Ours) 41.1 48 60.8 79
Code-T5 Distilled w/Augmentation (Ours) 41.9 68 61.1 80
GPT-4o-Mini (Auto-context) (Ours) 43.1 85 60.3 93
Table 8: Answer and program accuracy of different methods on GQA and VQAv2. With augmentation, distilled programs can achieve similar or better answer accuracy as prompting methods and improve program accuracy compared to not using augmentation.
The widely followed ViperGPT Surís et al. (2023) approach is to use a prompt containing an API and a set of human generated in-context examples. We use two models to evaluate prompting GPT-4o-Mini and Llama-3.2-1B Dubey et al. (2024), a distilled version of Llama 3. 25 human generated in-context examples are used per dataset for each model. In Table 8, we refer to these baselines as GPT-4o-Mini (w/human examples) and Llama-3.2-1B (w/human examples). From our method, we evaluate on two distilled CodeT5 models with and without augmentation. The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2). We also include the auto-context teacher model, notated as GPT-4o-Mini (Auto-context).
Table 9: Common errors and examples used during qualitative evaluation.
Error Question Program Explanation
Not executable Is the chair on the left or right side? chair = image_patch.find(‘chair’) side = choose_relationship(chair, image_patch, ‘left or right’) choose_relationship requires a list as input
API Violation What color is the running dog on the left? left = crop_position(‘running left’, image_patch) dog = left.find(‘dog’) color = dog.classify(‘color’) ‘Running left’ is not a direction or preposition
Contradicts Question What color is the car above the road? road = image_patch.find(‘road’) below_road = crop_position(‘below’, road) car = below_road.find(‘car’) car_color = car.classify(‘color’) The question states that the car is above the road, not below.
Does Not Answer Question Are there two tables? tables = image_patch.find(‘table’) num_tables = count(tables) The question asks if there are two tables, not how many tables there are.
Does not include all question information Is the blue toy small? toy = image_patch.find(‘toy’) is_small = toy.verify_property(‘small’) answer = is_small The question specifically asks about the blue toy.
We evaluate using two metrics: answer accuracy and program accuracy. Program accuracy involves a human (the authors) manually evaluating each program (without execution) for correctness. We randomly sample 100 questions from each dataset for evaluation. The same 100 questions are used to evaluate each method.
There are multiple ways to determine if a program is correct. We generally assume a program is correct unless it violates one of these criteria:
  • Not Executable: The program must be executable and return the correct data type (a string for VQA datasets).
  • API Violation: Visual programming APIs are designed to follow basic visual knowledge and reasoning. For example, the ‘find’ function is used with nouns while ‘verify_property’ is generally used for attributes. Clear violations such as trying to find an attribute (e.g. ‘find(green)’) or cropping with a verb (e.g. ‘crop_position(running)’) are considered incorrect.
  • Contradicts Question: Programs that assume a statement that directly conflicts with a statement in the question. For example, assuming an object is on the left, when the question states it is on the right.
  • Does Not Answer Question Programs that do not answer the question, even if the program correctly follows the API, are incorrect. Common examples are returning yes/no instead of choosing between two options such as left or right.
  • Does Not Include Vital Information From the Question If the question includes details about an object, then those details must be in the program.
Some examples can be seen in Table 9.
We can draw several observations from the results in Table 8. From the answer accuracy results, we see that with a small amount of data and a strong enough teacher, student models can achieve similar performance to the auto-context trained teacher. Auto-context generation also improves the teacher performance. The performance from prompting Llama is particularly low, consistent with previous findings showing that it is difficult to transfer prompts across LLMs Sclar et al. (2024) of similar size, let alone much smaller ones.
The program accuracy results are bit more surprising. For all of the methods, program accuracy is higher than answer accuracy especially for GPT-4o-Mini (Auto-context) which has 85% vs 43% and 93% vs 60% accuracy for GQA and VQAv2 (respectively). Data augmentation has a larger effect on program accuracy than question accuracy but a substantial gap still remains between the student and teacher models, indicating more room for improvement. Across the two datasets, we see a much higher overall program accuracy for VQAv2 than for GQA. Both datasets have fairly easy questions but VQAv2 has many that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count). Taken together, our evaluation suggests that the quality of the visual programs is much better than what answer accuracy alone would imply. We include additional analysis of program errors in Appendix A.
5 Conclusion
Our experiments show with template-based visual program distillation, only a small number of answer annotations are needed to achieve similar performance to prompt-based methods. Auto-context generation removes the burden of human generated program annotations while still retaining the same performance. Template-based augmentation improves both student/teacher prediction agreement and program accuracy compared to non-augmented approaches. The distilled student models achieve similar answer accuracy as the teacher at a fraction of the time and cost.
In conclusion, we show that the visual program ability of LLMs can be distilled into much smaller models with only a few question/answer pairs and no human-generated programs. For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 770-M coding model can become a visual program generator. We anticipate that use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their applications.
6 Limitations
There are several limitations to our work:
Indirect Evaluation Answer accuracy is an indirect evaluation of program correctness. As indicated by our program accuracy results shown in Section 4, answer-based evaluation is only a loose indicator of visual program correctness. Promising areas for future work include automatic program accuracy evaluation and closing the gap between answer accuracy and program accuracy.
Reliance on teacher model and API Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt. A future challenge is how to learn from a weaker teacher and/or unreliable API.
Teacher Data Efficiency Following prior works, if a program was incorrect, it was not used. However, LLMs have the ability to self-correct given the appropriate feedback. In settings where question/answer pairs are limited, such an approach could be more cost-effective than discarding the examples. Areas for future work include incorporating self-correction methods or using program correction models such as VDebugger Wu et al. (2024).
Program Execution Time Our method decreases program generation time, but not program execution time. With a single GPU, program execution time on average was 3.37 s and varies quite widely. The long execution time remains a bottleneck in different applications of visual programming.
Limited Program Complexity Existing VQA datasets are fairly simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM. Most real world applications are multi-step and would require more complex reasoning and knowledge skills than in the evaluation datasets.
Acknowledgement
This research is supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, ONR N00014-23-1-2383 and ONR N00014-21-1-2705. This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS230398 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program Boerner et al. (2023), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Program Analysis
Program Errors
In Figure 3, we show the frequency of different error types for each method on the GQA dataset. Note that incorrect programs can fall into multiple categories but in this classification, each incorrect program was counted only once. There are very few execution errors and most errors come from either contradicting the question or not answering it. Many of the Llama generated programs returned yes/no even if the question asked for a different string. Augmentation largely reduces such errors but also generates new errors.
Figure 3: The frequency of errors across the different categories for GQA program evaluation. Augmentation reduces the number of ‘Does Not Answer Question Mistakes.’
Qualitative Analysis
In Figure 4, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: student model trained without augmentation, with augmentation and the auto-context teacher. Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for the worse performing model, Llama-3.2-1B. In the second column, we see an example where the model trained without augmentation leaves out details mentioned in the question but the augmented model generates the correct program. All of the programs are incorrect in the last column but for different reasons. Both of the student models use the same program (apart from variable names) and make two mistakes. First the question asks about ‘not warm’ instead of ‘warm’ and the second is that the answer should be an object, not yes/no. The teacher program returns an object but still fails to recognize that the object should be ‘not warm’ even though the variable name includes ‘not warm’ in it. Questions involving negative properties are almost always missed by the teacher and student models.
Figure 4: 3 question/programs using no augmentation, augmentation and auto-context teacher. Simple comparison questions (left hand side) are almost always correct while questions with negations are almost always incorrect across the different methods.
Appendix B Training and Model Details
We used the following models for executing programs:
  1. 1.
    CLIP ViT-L/14 Radford et al. (2021)
  2. 2.
    InstructBLIP Flan-T5 XL Dai et al. (2023)
  3. 3.
    OWLv2 Base Patch 16 Ensemble Minderer et al. (2024)
Program generation settings for GPT can be found in Table 10. Template-based and direct training hyper-parameters can be found in Table 11. For CodeT5, the most important hyper-parameters were the learning rate and LoRA dropout rate. Training stopped when the training loss stopped decreasing.
All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were measured on an A40.
Setting Value
Temperature 0
Top_p 1.0
Frequency Penalty 0.0
Presence Penalty 0.0
Max Output Tokens 256
Table 10: GPT-4o-mini generation settings
Hyper-parameters Value
LoRA target modules All linear layers
LoRA rank 8
LoRA alpha 16
LoRA bias None
LoRA dropout 0.05 (no augmentation), 0.1 (augmentation)
LR 2e-4
Batch Size 16
Max Output Tokens 256
Table 11: Training and evaluation settings for CodeT5
Appendix C Changes to ViperGPT API
The following are major modifications made to the ViperGPT API Surís et al. (2023).
Figure 5: Difference in program annotations when a VLM is used as a backup model for the question ‘Is the ground blue or brown?’ The highlighted portion is removed from all program annotations used.
  1. 1.
    Program annotations were modified not to use a vision-language model (VLM) when the program fails (see Figure 5 for an example). In the original ViperGPT API, examples in the API included a line to directly query a VLM if other parts of the program failed such as when no object is found. The performance using the original ViperGPT code decreases considerably when the VLM backup lines are removed from the API as shown in Table 12.
  2. 2.
    An object is always returned by the object detector.
  3. 3.
    Program annotations did not include parts of the program that were shared among all examples.
  4. 4.
    Several new functions were added.
    1. (a)
      Verify Relationship: Given two objects and a relation, return a boolean whether the objects satisfy that relationship.
    2. (b)
      Choose Relationship: Given two objects, return the relationship between the two.
    3. (c)
      Crop Position: Crop part of the image based on a position relative to an object
Use of VLM Backup GQA-Test Dev
ViperGPT with VLM Backup 47.3
ViperGPT without VLM Backup 26.0
Table 12: Change in GQA test-dev accuracy using original ViperGPT API when not using a VLM when the program fails
Appendix D Prompt
Instructions
For each question provided, generate a Python program that includes a return statement. Assume that image_patch = ImagePatch(image) is already defined. The final output of the program should always be a string.
ImagePatch
Attributes
1. **cropped_image**
Type: array
Description: An array representing the cropped image.
2. **left**
Type: int
Description: The left border of the crop’s bounding box.
3. **lower**
Type: int
Description: The bottom border of the crop’s bounding box.
4. **right**
Type: int
Description: The right border of the crop’s bounding box.
5. **upper**
Type: int
Description: The top border of the crop’s bounding box.
Methods
1. **find(object_name: str) -> List[ImagePatch]**
Description: Returns a list of image patches containing the specified object.
Notes: find should not be the last operation in a program.
Examples:
image_patch.find(’chair’)
image_patch.find(’table’)
2. **crop_position(direction: str, reference_patch: ImagePatch) -> ImagePatch**
Description: Returns a new image patch in the specified direction relative to the reference_patch. Directions can include ’left’, ’right’, ’above’, ’below’, ’on’, ’in front’, etc.
Notes: The result of crop_position should not be immediately indexed on the next line. The second argument is always the original image_patch.
Examples:
image_patch.crop_position(’left’, image_patch)
image_patch.crop_position(’above’, image_patch)
3. **verify_property(property_name: str) -> bool**
Description: Returns True if the object contains the specified property; otherwise, False.
Notes: Can only be called on an image patch.
Examples:
image_patch.verify_property(’red’)
image_patch.verify_property(’running’)
4. **classify(options: Union[str, List[str]]) -> str**
Description: Given a category (e.g., ’color’, ’material’, ’furniture’) or a list of options, returns the best option for the image patch.
Notes: The input should not be ’object’.
Examples:
image_patch.classify([’red’, ’blue’])
image_patch.classify(’color’)
5. **simple_query(question: str) -> str**
Description: Answers questions about the image, especially ambiguous ones (e.g., ’Who is riding?’).
Examples:
image_patch.simple_query(’Who is riding?’)
General Functions
1. **filter_img(image_patches: List[ImagePatch], criteria: str) -> List[ImagePatch]**
Description: Filters the list of image patches based on the given criteria. The criteria can be an action, attribute, or object.
Examples:
filter_img(image_patches, ’red’)
filter_img(image_patches, ’running’)
2. **choose_relationship(patch1: Union[ImagePatch, List[ImagePatch]], patch2: Union[ImagePatch, List[ImagePatch]], relationships: Union[List[str], str]) -> str**
Description: Chooses the relationship that best matches the two patches from the provided options.
Examples:
choose_relationship(image_patch1, image_patch2, [’on top of’, ’next to’])
choose_relationship(image_patch1, image_patch2, [’left’, ’right’])
3. **verify_relationship(patch1: Union[ImagePatch, List[ImagePatch]], patch2: Union[ImagePatch, List[ImagePatch]], relationship: str) -> str**
Description: Returns ’yes’ or ’no’ based on whether the specified relationship holds between the two patches.
Examples:
verify_relationship(image_patch1, image_patch2, ’on top of’)
verify_relationship(image_patch1, image_patch2, ’left’)
4. **exists(patches: Union[ImagePatch, List[ImagePatch]]) -> bool**
Description: Checks whether any of the provided image patches exist.
Notes: If used as the last operation, it should be followed by bool_to_yesno().
Examples:
exists(image_patches)
5. **bool_to_yesno(value: bool) -> str**
Description: Converts a boolean value to ’yes’ or ’no’. Used to convert outputs of verify_property and exists.
Examples:
bool_to_yesno(exists(image_patches))
Here are some examples of how to write programs:
{examples}
Additional Notes
- You may utilize standard Python functions within your programs.
- Do not include comments.
- Only return the program.
- Do not define the function.
- Functions never return None.
- The last line of each program should be answer =
Appendix E Variable Renamer
class VariableRenamer(ast.NodeTransformer):
def __init__(self, skip_vars=None):
self.counter = 1 # For general variables (var1, var2, ...)
self.temp_counter = 1 # For comprehension/loop variables (temp_var_1, ...)
self.name_map = {}
self.skip_vars = set(skip_vars) if skip_vars else set()
def _new_name(self):
name = f"var{self.counter}"
self.counter += 1
return name
def _new_temp_name(self):
[…]
self.temp_counter += 1
return name
def rename_target(self, target):
"""Rename normal assignment or loop targets, skipping those in skip_vars."""
if isinstance(target, ast.Name):
if target.id in self.skip_vars:
return target
if target.id not in self.name_map:
self.name_map[target.id] = self._new_name()
target.id = self.name_map[target.id]
elif isinstance(target, (ast.Tuple, ast.List)):
for elt in target.elts:
self.rename_target(elt)
return target
def visit_Name(self, node):
if isinstance(node.ctx, (ast.Store, ast.Load, ast.Del)):
if node.id in self.skip_vars:
return node
if node.id in self.name_map:
node.id = self.name_map[node.id]
return node
def visit_Assign(self, node):
node.value = self.visit(node.value)
node.targets = [self.rename_target(t) for t in node.targets]
return node
def rename_within(self, node, old_name, new_name):
"""Recursively replace occurrences of old_name with new_name within the node."""
class NameReplacer(ast.NodeTransformer):
def visit_Name(self, n):
if n.id == old_name:
n.id = new_name
return n
replacer = NameReplacer()
return replacer.visit(node)
def visit_For(self, node):
# Enhanced handling for For loops to propagate renaming within the loop body.
if isinstance(node.target, ast.Name) and node.target.id not in self.skip_vars:
old_name = node.target.id
new_temp = self._new_temp_name()
node.target.id = new_temp
# Visit and rename within ’iter’, ’body’, and ’orelse’
node.iter = self.visit(node.iter)
node.body = [self.rename_within(self.visit(n), old_name, new_temp) for n in node.body]
if node.orelse:
node.orelse = [self.rename_within(self.visit(n), old_name, new_temp) for n in node.orelse]
else:
node.target = self.rename_target(node.target)
node.iter = self.visit(node.iter)
node.body = [self.visit(n) for n in node.body]
if node.orelse:
node.orelse = [self.visit(n) for n in node.orelse]
return node
def visit_While(self, node):
node.test = self.visit(node.test)
node.body = [self.visit(n) for n in node.body]
if node.orelse:
node.orelse = [self.visit(n) for n in node.orelse]
return node
def visit_ListComp(self, node):
for gen in node.generators:
if isinstance(gen.target, ast.Name) and gen.target.id not in self.skip_vars:
old_name = gen.target.id
new_temp = self._new_temp_name()
gen.target.id = new_temp
node.elt = self.rename_within(node.elt, old_name, new_temp)
gen.ifs = [self.rename_within(if_clause, old_name, new_temp) for if_clause in gen.ifs]
for inner_gen in node.generators:
inner_gen.target = self.rename_within(inner_gen.target, old_name, new_temp)
else:
gen.target = self.rename_target(gen.target)
gen.iter = self.visit(gen.iter)
node.elt = self.visit(node.elt)
for gen in node.generators:
gen.ifs = [self.visit(if_clause) for if_clause in gen.ifs]
return node
def visit_GeneratorExp(self, node):
for gen in node.generators:
if isinstance(gen.target, ast.Name) and gen.target.id not in self.skip_vars:
old_name = gen.target.id
new_temp = self._new_temp_name()
gen.target.id = new_temp
node.elt = self.rename_within(node.elt, old_name, new_temp)
gen.ifs = [self.rename_within(if_clause, old_name, new_temp) for if_clause in gen.ifs]
for inner_gen in node.generators:
inner_gen.target = self.rename_within(inner_gen.target, old_name, new_temp)
else:
gen.target = self.rename_target(gen.target)
gen.iter = self.visit(gen.iter)
node.elt = self.visit(node.elt)
for gen in node.generators:
gen.ifs = [self.visit(if_clause) for if_clause in gen.ifs]
return node
def visit_With(self, node):
for item in node.items:
if item.optional_vars and isinstance(item.optional_vars, ast.Name) and item.optional_vars.id not in self.skip_vars:
item.optional_vars.id = self._new_temp_name()
elif item.optional_vars:
item.optional_vars = self.rename_target(item.optional_vars)
item.context_expr = self.visit(item.context_expr)
node.body = [self.visit(n) for n in node.body]
return node
# Additional visitor methods for other constructs can be added here.
def format_assignments(source_code: str) -> str:
Remove spaces around the equals sign in single-line assignment statements
without altering multi-line assignments.
This function ensures that:
- Single-line assignments have no spaces around ’=’.
- Multi-line assignments are left intact to preserve code correctness.
[…]
formatted_lines = []
[…]
# Track the balance of parentheses, brackets, and braces
paren_balance = 0
for line in lines:
stripped_line = line.strip()
# Update paren_balance
paren_balance += line.count(’(’) - line.count(’)’)
paren_balance += line.count(’[’) - line.count(’]’)
paren_balance += line.count(’{’) - line.count(’}’)
# If paren_balance > 0, we’re inside a multi-line expression
if paren_balance > 0:
formatted_lines.append(line)
continue
# Attempt to match an assignment statement
match = assignment_pattern.match(line)
if match:
indent, var, expr = match.groups()
# Remove spaces around ’=’ and reconstruct the line
formatted_line = f"{indent}{var}={expr}"
formatted_lines.append(formatted_line)
else:
# Non-assignment lines are added directly
formatted_lines.append(line)
# Join the lines back into a single string
[…]
def replace_variables(code: str, convert_to_source: bool = True) -> Union[str, ast.AST]:
skip_list = {"image_patch", "answer"} # Variables not to rename
tree = ast.parse(code)
renamer = VariableRenamer(skip_vars=skip_list)
new_tree = renamer.visit(tree)
ast.fix_missing_locations(new_tree)
new_source = ast.unparse(new_tree)
formatted_source = format_assignments(new_source)
return formatted_source
Given a generated program we call the function ‘replace_variables’ which uses and abstract-syntax tree to rename variables both in and outside different types of loops.
Appendix F Variable Replacement
Algorithm 1 Argument Replacement
  Extract arguments per function
  for func in functions do
     for arg in arguments do
        if arg in category then
           Random sample from category
        else
           Random sample from generic object list
        end if
     end for
  end for
The general algorithm for replacing an argument in a program can be seen in Algorithm 1. For each named function or method in the API, we extract the arguments. For GQA, if the argument is already in a pre-defined catgory, we randomly sample from that category. Otherwise we randomly sample an object. Some example categories and options can be seen in Table 13. The process for VQA is similar except there is no pre-defined list. Instead we mask out the argument in the question and generate a replacement using BERT or BART if the argument is a phrase. For full questions, we use word tokenization (default NLTK Tokenization) and POS-tagging with NLTK Bird et al. (2009) to determine where to place masks. We randomly sample from the top-50 results. Both the BERT and BART models are large uncased with 340 and 406 M parameters.
We perform the process above if an argument is selected for replacement during training. Each argument has a probability $p=0.5$ to be selected.
Table 13: Argument Categories and Options
Category Name Category Examples
Color red, blue, green, yellow, purple, black, white, orange, pink, brown, gray, indigo, cyan magenta, tan, silver
Activities running, walking, snowboarding, flying, splashing, tossing, riding, standing, hugging hanging, breaking, pulling, decorating, facing, preparing pouring, pointing, laughing
Relation picking up, in front of behind, above, below next to, near, far away close , following, on top, beside, walking on, attached, left, right, diagonal